Grade 12General Physics

capacity of vessel is 5700m3. the vessel is filled with water. suppose that it takes 12 hours to drain vessel.what is the mass flow rat of water from the vessel? density of water is 1g/cm3

BHAVANA RM
11 Years agoGrade 12

2 Answers

Abhishek
ApprovedApproved Tutor Answer11 Years ago
it will be 131.944 kg/sec
Nikhil KUMAR Choudhary
7 Years ago
The density of water is 1000kg/m^3
Volume of water is 5700m^3
Time to get empty =12hour
Per hour is =5700/12
Per sec is =5700/12*3600m^3
Density =mass *volume =1000*5700/12*3600=131.9kg/s
Density =mass/volume